OpenStudy (anonymous):

How to solve it? √(2k+1) -26=-5 thx.

OpenStudy (anonymous):

220

OpenStudy (agreene):

add 26 to both sides, then square both sides. solve from there.

OpenStudy (chriss):

sqrt(2k+1)-26 = -5 +26 +26 sqrt(2k+1) = 21 2k+1=441 2k=440 k=220